OpenStudy (anonymous):

what is the equation of a line that passes through (7,8) and has a slope of -3 ?

OpenStudy (ilovecake):

http://www.enotes.com/homework-help/what-equation-line-passing-through-point-0-8-459325 Here is something close to your question

OpenStudy (gorv):

standard equation y=mx+c m=slope=-3 y=-3x+c for finding value of c put value of x and y 8=-3*7+c 8=-21+c c=29 so y=-3x+29 is your equation

OpenStudy (ilovecake):

@itsgermanbruh